OpenStudy (anonymous):

identify the transformation of f(x) that each function represents 1. g(x) = f(x - 5) 2. g(x) = 3f(x)

OpenStudy (anonymous):

and can you show me the work so i can learn this

jimthompson5910 (jim_thompson5910):

In general, if you replace x with x-h, you are shifting the entire graph of f(x) `h` units to the left. That is if h is positive if h is negative, then you go the same number of units but to the right

OpenStudy (anonymous):

i don't have h do you mean g?

jimthompson5910 (jim_thompson5910):

notice how f(x-5) and f(x-h) match up

OpenStudy (anonymous):

because your replacing 5 with h?

jimthompson5910 (jim_thompson5910):

yeah so in your specific case, h = 5 so when we go from f(x) to f(x-5) we shift all of f(x) 5 units to the right. Sorry I mixed up the order above. I mixed up left and right, my bad

jimthompson5910 (jim_thompson5910):

h is positive, you shift right h is negative, you shift left

OpenStudy (anonymous):

oh okay so thats all i write?

OpenStudy (anonymous):

and for the second one what dould it be since it's multiplication

jimthompson5910 (jim_thompson5910):

`oh okay so thats all i write?` yeah all you need to say really is that the entire graph shifts 5 units to the right

jimthompson5910 (jim_thompson5910):

that 3 out front means every y coordinate gets tripled visually what that does is vertically stretch the graph by a factor of 3. it's basically 3 times taller than it was before

jimthompson5910 (jim_thompson5910):

example: say we know the point (2,5) is on f(x) we know for a fact that (2,15) is on g(x) because 3*f(2) = 3*5 = 15

jimthompson5910 (jim_thompson5910):

imagine that happens with every point on f(x)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i have more do you mind helping me i've missed school cause i got really sick and this is make up work but i never learnt this

jimthompson5910 (jim_thompson5910):

I'll help with 2 more

OpenStudy (anonymous):

i have a whole bunch thought...

OpenStudy (anonymous):

do you mind if we schedule a time tomorrow so you can help me?